Christia is standing on the top of the head of John's shadow. If she is 5 feet tall and his height is 6 feet, and the length of their shadow is 15 feet, find the length of Christia's shadow.
Answer by stanbon(75887)

Christia is standing on the top of the head of John's shadow. If she is 5 feet tall and his height is 6 feet, and the length of their shadow is 15 feet, find the length of Christia's shadow.
------
15/(6+5) = s/5
-----
shadow = 75/11 = 6.82 ft
=============
Cheers,
Stan H.
--------------